you invested 4000 in 2 stocks. one is 5% and the other 7% annual interest. At the end of the year total interest was 230. how much was invested at each rate?
Answer by texttutoring(324)

Let x=amount invested at 5%
Let y=amount invested at 7%

Equation 1: x+y=4000
Equation 2: 0.05x + 0.07y=230

Isolate for y in Eqn. 1 and substitute it into Eqn 2:

y=4000-x

Put this value of y into Eqn 2 instead of y:

0.05x+0.07(4000-x) = 230
0.05x +280-0.07x = 230
-0.02x = -50
x=2500

Plug this value of x into Eqn 1 to solve for y:

x+y=4000
y=4000-x
y=4000-2500
y=1500

So $2500 was invested at 5%, and $1500 was invested at 7%.